OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 − 6x + 4

terenzreignz (terenzreignz):

You can take this to faith, but... when you have a quadratic equation of the form \[\huge ax^2+bx+c\]Then its vertex is at the point \[\huge \left(-\frac{b}{2a} \ \ , \ \ \frac{4ac-b^2}{4a}\right)\] Or, you could just get this... \[\Large -\frac{b}{2a}\]And evaluate your equation with x equal to this... that'd be the y-value of the vertex Have fun, and welcome to Openstudy :D